OpenStudy (anonymous):

Solve the equation 2/7=6x/4x+17. The solution of the equation is?

OpenStudy (astrophysics):

Is this your question \[\frac{ 2 }{ 7 } = \frac{ 6 }{ 4 }x+17\]

OpenStudy (astrophysics):

Or is it \[\frac{ 2 }{ 7 } = \frac{ 6x }{ 4x+17 }\]

OpenStudy (anonymous):

the second one

OpenStudy (astrophysics):

Ok well, all you have to do is cross multiply, do you know how?

OpenStudy (anonymous):

no

OpenStudy (astrophysics):

 now simplify and solve for x